# Object Segmentation-Assisted Inter Prediction for Versatile Video Coding

### v1

@@ VI Segmentation-Assisted Rate-Distortion Optimization for Motion Estimation and
In codec, the SAIP mode is added as a regular prediction mode in the mode decision list for each coding unit. For a coding block (CB) to try the SAIP mode, first, the primary candidate list is established. To provide the various partition candidates for accurate partition estimation, the primary candidate list is designed to be longer than the secondary candidate list. Each primary candidate contains MV, Ref and Partition with index . Meanwhile, each primary candidate indicates a different reference block, and the different segmentation mask (representing the different regions by 1/0-value) can be translated to suppose the different motion area partition of current coding block (Partition). For the motion representation of each MV in the coding block, it is hard to determine whether the MV represents the 1-value region or the 0-value region’s motion. Therefore, the reverse index is added to identify the distribution of primary region in the segmentation mask. When the reverse index is 0, the 1-value indicates the primary region and the 0-value indicates the other parts. When the reverse index is 1, the opposite is true. Once the partition and primary region are fixed, the other region corresponds to the secondary region. Second, the secondary candidate list is constructed according to SA-CD and Partition, the MV and Ref with index can be selected to represent the motion of other parts in the coding block.
@@
With the primary and secondary candidates constructed, the best combination of these candidates is determined by two stages.
VI-1 Stage 1
@@
The combined primary and secondary candidates are used to conduct the SA-MC to generate the prediction of the coding block. The sum of the absolute transformed differences (SATD) of luma between the prediction and the original signal is computed as . The rate-distortion (RD) cost for , and can be sorted by
@@ VI-2 Stage 2
For each coding block, 70 (the length of primary candidate list) 7 (the length of secondary candidate list) 2 (reverse test) = 980 times are calculated and compared in Stage 1 of SA-RDO, and 4 (the length of combined candidate list) 2 (residual coding, use or skip) = 8 times are calculated and compared in Stage 2 of SA-RDO, 988 times maximum calculation in total.
@@
| Class | Sequence | Low-delay P (%) | Low-delay B (%) | Random Access (%) | ||||||
| Name | Y | U | V | Y | U | V | Y | U | V | |
| ClassA1 (3840x2160) | Tango2 | - | - | - | - | - | - | -0.21% | 0.21% | -0.11% |
| FoodMarket4 | - | - | - | - | - | - | -0.25% | -0.25% | -0.20% | |
| Campfire | - | - | - | - | - | - | -0.16% | 0.02% | -0.12% | |
| ClassA2 (3840x2160) | CatRobot1 | - | - | - | - | - | - | -0.69% | -0.01% | -0.37% |
| DaylightRoad2 | - | - | - | - | - | - | -0.21% | -0.37% | -0.38% | |
| ParkRunning3 | - | - | - | - | - | - | -0.06% | -0.01% | 0.03% | |
| ClassB (1920x1080) | MarketPlace | -0.43% | 0.31% | -0.62% | -0.46% | -0.78% | -0.99% | -0.28% | -0.13% | -0.07% |
| RitualDance | -0.23% | -0.35% | -0.48% | -0.28% | 0.42% | 0.17% | -0.18% | 0.27% | -0.12% | |
| Cactus | -0.28% | -0.47% | -0.38% | -0.28% | -0.24% | -0.19% | -0.30% | 0.33% | -0.40% | |
| BasketballDrive | -0.29% | -0.43% | -0.66% | -0.25% | -0.28% | -0.33% | -0.20% | 0.75% | -0.21% | |
| BQTerrace | -0.03% | -0.09% | -0.11% | -0.19% | 0.47% | 0.49% | -0.34% | -0.19% | -0.40% | |
| ClassC (832x480) | BasketballDrill | -1.83% | -1.03% | -1.83% | -1.14% | -0.74% | -1.00% | -0.79% | -0.08% | -0.41% |
| BQMall | -1.98% | -2.84% | -2.41% | -0.84% | -1.07% | -0.63% | -0.66% | -1.08% | -0.70% | |
| PartyScene | -0.27% | -0.61% | -1.03% | -0.17% | -0.36% | -0.16% | -0.17% | -0.27% | 0.03% | |
| RaceHorsesC | -0.95% | -1.23% | -0.39% | -0.23% | -0.13% | 0.30% | -0.37% | -0.40% | -0.06% | |
| ClassE (1280x720) | FourPeople | -0.93% | -0.64% | -0.17% | -0.46% | 0.29% | -0.36% | -0.41% | -0.35% | -0.21% |
| Johnny | -1.89% | -1.29% | -1.30% | -1.12% | -2.02% | 0.46% | -0.67% | -0.25% | -0.34% | |
| KristenAndSara | -0.68% | -1.63% | 0.02% | -0.38% | -0.80% | 0.66% | -0.48% | -0.77% | -0.18% | |
| Overall | -0.82% | -0.86% | -0.78% | -0.49% | -0.45% | -0.13% | -0.37% | -0.14% | -0.23% | |
| ClassD (416x240) | BasketballPass | -0.61% | -0.38% | -1.17% | -0.48% | -0.97% | -0.74% | -0.44% | 0.55% | -0.21% |
| BQSquare | -0.78% | -2.28% | 3.65% | -0.42% | -1.73% | 1.76% | -0.53% | -0.60% | -0.37% | |
| BlowingBubbles | -1.02% | -1.17% | -1.98% | -0.63% | -0.16% | 0.39% | -0.50% | -0.81% | -1.06% | |
| RaceHorses | -0.83% | -0.21% | -0.90% | -0.10% | 0.88% | 0.15% | -0.34% | 0.94% | 0.49% | |
| Class | Low-delay P | Low-delay B | Random Access | |||
| EncT | DecT | EncT | DecT | EncT | DecT | |
| A | - | - | - | - | 183% | 116% |
| B | 179% | 111% | 171% | 112% | 180% | 112% |
| C | 159% | 109% | 147% | 113% | 154% | 115% |
| E | 178% | 113% | 148% | 120% | 179% | 118% |
| Overall | 172% | 111% | 155% | 115% | 174% | 115% |
| D | 135% | 108% | 123% | 119% | 116% | 117% |
VII Experiments
VII-A Experimental Settings
VII-A1 Learned Segmentation Inference
@@
VII-A2 Encoding Configurations
To evaluate the performance of SAIP in VVC, SAIP is implemented into the VVC reference software VTM-12.0 222https://vcgit.hhi.fraunhofer.de/jvet/VVCSoftware_VTM. Three main profile configurations (Low-delay P (LDP), Low-delay B (LDB), Random Access (RA)) and test sequences specified in [42] are used as the test conditions. Following the common test conditions (CTC)[42], four quantization parameters are tested: 22, 27, 32 and 37. For different quantization parameters, different segmentation configurations of the SCNet and STCN are chosen. Finally, the Bjontegaard Delta bit-rate (BD-rate) [43] is used as the objective metric to evaluate the coding performance.
VII-B Performance
@@
VII-B1 Overall Performance Under Common Test Conditions
The R-D performance of the entire segmentation-assisted inter prediction framework on the VVC common test sequences is illustrated in Table I . Y, U, and V represent the R-D performance gain of the three channels of YUV. We can see that our proposed SAIP can achieve on average, 0.82%, 0.49% and 0.37% (marked by font bold), and achieve up to 1.98%, 1.14% and 0.79% (marked by underline) BD-rate reduction (Y component) on VTM-12.0 for all sequences under the LDP, LDB, and RA configurations. The experimental results show that the proposed framework performs better for sequences with more moving objects and complex motion fields, such as BasketballDrill, BQMall, CatRobot, and Marketplace.
| Low-delay B Configuration (%) | |||||||||||||||||||||||
@@ VII-B1 Overall Performance Under Common Test Conditions
| Name | Y | Y | U | V | Y | U | V | Y | U | V | Y | U | V | ||||||||||
| MarketPlace | - | - | - | - | -16.81% | -12.53% | -20.41% | -4.86% | 20.24% | 22.04% | -2.38% | -2.09% | -2.38% | ||||||||||
| RitualDance | - | - | - | - | -25.05% | -20.87% | -20.66% | -8.90% | 17.14% | 16.08% | -2.15% | -1.98% | -1.91% | ||||||||||
| Cactus | - | -1.00% | -0.90% | -1.00% | -20.36% | -6.45% | -6.35% | -9.88% | 22.29% | 22.37% | -0.80% | -0.77% | -0.86% | ||||||||||
| BasketballDrive | -0.08% | -0.70% | -0.80% | -1.10% | -24.02% | -21.09% | -21.25% | -9.47% | 27.71% | 23.75% | -2.20% | -3.00% | -2.88% | ||||||||||
| BQTerrace | - | -1.60% | -2.00% | -1.70% | -16.10% | -15.34% | -28.19% | -8.08% | 18.61% | 11.56% | -1.48% | -1.01% | -0.94% | ||||||||||
| Kimono | -0.14% | -0.90% | -1.00% | -1.40% | -15.01% | 4.40% | 0.68% | -6.15% | 20.14% | 20.14% | -1.17% | -1.00% | -0.87% | ||||||||||
| ParkScene | - | -0.60% | -0.40% | -0.70% | -16.60% | -3.55% | -4.14% | -6.89% | 20.26% | 22.02% | -0.23% | -0.63% | -0.55% | ||||||||||
| BasketballDrill | - | -1.30% | -1.50% | -1.30% | -29.14% | -15.11% | -14.63% | -19.76% | 9.37% | 10.45% | -4.34% | -4.27% | -4.91% | ||||||||||
| BQMall | - | -1.50% | -1.10% | -1.00% | -22.74% | -14.37% | -13.82% | -12.03% | 14.86% | 14.43% | -4.47% | -5.91% | -5.14% | ||||||||||
| PartyScene | - | -1.80% | -1.40% | -1.10% | -16.82% | 0.69% | -1.07% | -9.07% | 21.47% | 22.33% | -1.16% | -1.10% | -1.37% | ||||||||||
| RaceHorsesC | -0.95% | -0.60% | -0.90% | -0.80% | -12.52% | 10.91% | 10.28% | -5.06% | 30.48% | 34.81% | -2.21% | -2.97% | -3.63% | ||||||||||
| BasketballPass | - | -0.80% | -0.80% | -1.00% | -17.88% | -10.36% | -6.16% | -7.69% | 23.35% | 24.07% | -2.41% | -3.60% | -4.02% | ||||||||||
| BQSquare | - | -1.90% | -1.60% | -2.00% | -13.55% | -10.62% | -12.38% | -8.34% | 7.68% | 5.87% | -1.26% | 0.17% | 1.14% | ||||||||||
| BlowingBubbles | -0.30% | -1.70% | -1.40% | -1.60% | -12.62% | 6.62% | 2.90% | -5.55% | 24.08% | 24.18% | -2.18% | -2.35% | -1.14% | ||||||||||
| RaceHorses | - | -1.00% | -1.20% | -1.00% | -13.47% | 7.69% | 7.84% | -5.10% | 29.82% | 29.66% | -2.11% | -2.81% | -2.30% | ||||||||||
| FourPeople | - | -0.70% | -0.60% | -1.20% | -25.13% | -28.39% | -29.28% | -10.46% | 0.92% | -0.80% | -2.23% | -1.48% | -1.86% | ||||||||||
| Johnny | - | -0.90% | -1.00% | -1.20% | -25.81% | -30.38% | -34.35% | -10.25% | 9.35% | 6.11% | -3.80% | -3.04% | -2.73% | ||||||||||
| KristenAndSara | - | -1.10% | -1.30% | -0.90% | -23.97% | -26.32% | -28.35% | -8.89% | 6.64% | 5.79% | -1.53% | -1.07% | 2.21% | ||||||||||
| Average | -0.37% | -1.13% | -1.12% | -1.13% | -20.72% | -12.74% | -15.11% | -8.69% | 18.02% | 17.49% | -2.12% | -2.16% | -1.90% | ||||||||||
@@
In VVC reference software (VTM-12.0), the technology of inter prediction most relevant to SAIP is the geometric partitioning mode (GEO)[11]. Both ours and GEO aim to achieve a flexible description of the motion field and match the appropriate prediction for different regions in a block. In Table I , under the LDB and RA configurations, GEO is enabled in the anchor by default. In the comparison with anchor, SAIP achieves good performance with GEO enabled. It can be concluded that SAIP can efficiently deal with some complex motion scenarios that GEO is difficult to address.
VII-B2 Complexity
The encoding/decoding time complexity of our proposed method is illustrated in Table II . EncT and DecT represent the encoding/decoding time with the inference time of segmentation network, the network is inferred in parallel with the syntax parsing. The codec time is tested on CPU, and the model is Intel(R) Core(TM) i7-11700 @2.50GHz. The inference time of network is tested on GPU, the models are aligned with the [27, 33], and the computational complexity is 119 kMAC/pixel.
@@ VII-B2 Complexity
| Low-delay P Configuration (%) | Random Access Configuration (%) | |||||||||||||||||||||||||

### v2

@@ VII-A Rate-Distortion Optimization for Uni-directional Prediction
In codec, the SAIP mode is added as a regular prediction mode in the mode decision list for each coding unit. For a coding block (CB) to try the uni-directional SAIP mode, first, the uni-directional primary candidate list is established. To provide the various partition candidates for accurate partition estimation, the primary candidate list is designed to be longer than the secondary candidate list. Each primary candidate contains MVprimary,α, Refprimary,α and PartitionCB,α with index . Meanwhile, each primary candidate indicates a different reference block, and the different segmentation mask (representing the different regions by 1/0-value) can be translated to suppose the different motion area partition of current coding block (PartitionCB,α). For the motion representation of each MVprimary,α in the coding block, it is hard to determine whether the MVprimary,α represents the 1-value region or the 0-value region’s motion. Therefore, the reverse index is added to identify the distribution of primary region in the segmentation mask. When the reverse index is 0, the 1-value indicates the primary region and the 0-value indicates the other parts. When the reverse index is 1, the opposite is true. Once the partition and primary region are fixed, the other region corresponds to the secondary region. Second, the uni-directional secondary candidate list is constructed according to SA-CD and PartitionCB,α, the MVsecondary,β and Refsecondary,β with index can be selected to represent the motion of other parts in the coding block.
@@
With the uni-directional primary and secondary candidates constructed, the best combination of these candidates is determined by two stages.
VII-A1 Stage 1
@@
The combined uni-directional primary and secondary candidates are used to conduct the SA-MC to generate the prediction of the coding block. The sum of the absolute transformed differences (SATD) of luma between the prediction and the original signal is computed as . The rate-distortion (RD) cost for , and can be sorted by
@@ VII-A2 Stage 2
For the uni-directional prediction process of each coding block, 71 (the length of primary candidate list) 7 (the length of secondary candidate list) 2 (reverse test) = 994 times are calculated and compared in Stage 1 of SA-RDO, and 4 (the length of combined candidate list) 2 (residual coding, use or skip) = 8 times are calculated and compared in Stage 2 of SA-RDO, 1002 times maximum calculation in total.
VII-B Rate-Distortion Optimization for Bi-directional Prediction
In the bi-directional case of SA-RDO, each bi-directional primary candidate contains MVprimary,α,f, Refprimary,α,f, PartitionCB,α,f, and MVprimary,α,b, Refprimary,α,b, PartitionCB,α,b with index . The indicates the forward reference direction, and indicates the backward reference direction. For different reference directions, the motion information of the corresponding reference direction of primary candidate indicates a reference block (retrieved by MVprimary,α,f/Refprimary,α,f or MVprimary,α,b/Refprimary,α,b), and a segmentation mask (PartitionCB,α,f or PartitionCB,α,b) can be translated to suppose the different motion area partition of current coding block in the corresponding reference direction. With the assistance of partition candidate (PartitionCB,α,f and PartitionCB,α,b) in the corresponding reference direction, the subsequent prediction process of each reference direction is the same as the uni-directional case of SA-RDO, and each bi-directional prediction candidate is obtained by the weighted average of different reference direction’s prediction.
| Syntax | Descriptor |
| merge_data() { | |
| saip_flag | ae(v) |
| if( saip_flag ){ | |
| saip_merge_flag | ae(v) |
| if( saip_merge_flag ){ | |
| saip_merge_idx | ae(v) |
| else{ | |
| saip_mmvd_cand_idx | ae(v) |
| saip_mmvd_distance_idx | ae(v) |
| saip_mmvd_direction_idx | ae(v) |
| saip_back_idx | ae(v) |
| saip_reverse_idx | ae(v) |
VII-C Complexity Optimization
@@
| Class | Sequence | Low-delay P (%) | Low-delay B (%) | Random Access (%) | ||||||||||||
| Name | Y | U | V | Ratio | Y | U | V | Ratio | Y | U | V | Ratio | ||||
| ClassA1 (3840x2160) | Tango2 | - | - | - | - | - | - | - | - | -0.21% | 0.21% | -0.11% | 1.19% | |||
| FoodMarket4 | - | - | - | - | - | - | - | - | -0.25% | -0.25% | -0.20% | 1.63% | ||||
| Campfire | - | - | - | - | - | - | - | - | -0.16% | 0.02% | -0.12% | 1.52% | ||||
| Average | - | - | - | - | - | - | - | - | -0.20% | -0.01% | -0.14% | 1.45% | ||||
| ClassA2 (3840x2160) | CatRobot1 | - | - | - | - | - | - | - | - | -0.69% | -0.01% | -0.37% | 2.12% | |||
| DaylightRoad2 | - | - | - | - | - | - | - | - | -0.21% | -0.37% | -0.38% | 1.22% | ||||
| ParkRunning3 | - | - | - | - | - | - | - | - | -0.06% | -0.01% | 0.03% | 0.72% | ||||
| Average | - | - | - | - | - | - | - | - | -0.32% | -0.13% | -0.24% | 1.35% | ||||
| ClassB (1920x1080) | MarketPlace | -0.43% | 0.31% | -0.62% | 2.89% | -0.46% | -0.78% | -0.99% | 2.78% | -0.28% | -0.13% | -0.07% | 1.36% | |||
| RitualDance | -0.23% | -0.35% | -0.48% | 2.15% | -0.28% | 0.42% | 0.17% | 2.27% | -0.18% | 0.27% | -0.12% | 1.21% | ||||
| Cactus | -0.28% | -0.47% | -0.38% | 2.34% | -0.28% | -0.24% | -0.19% | 2.37% | -0.30% | 0.33% | -0.40% | 1.58% | ||||
| BasketballDrive | -0.29% | -0.43% | -0.66% | 2.52% | -0.25% | -0.28% | -0.33% | 1.69% | -0.20% | 0.75% | -0.21% | 1.45% | ||||
| BQTerrace | -0.03% | -0.09% | -0.11% | 0.73% | -0.19% | 0.47% | 0.49% | 1.68% | -0.34% | -0.19% | -0.40% | 1.11% | ||||
| Average | -0.25% | -0.21% | -0.45% | 2.13% | -0.30% | -0.08% | -0.17% | 2.16% | -0.26% | 0.21% | -0.24% | 1.34% | ||||
| ClassC (832x480) | BasketballDrill | -1.83% | -1.03% | -1.83% | 10.73% | -1.14% | -0.74% | -1.00% | 7.59% | -0.79% | -0.08% | -0.41% | 4.45% | |||
| BQMall | -1.98% | -2.84% | -2.41% | 10.39% | -0.84% | -1.07% | -0.63% | 6.54% | -0.66% | -1.08% | -0.70% | 4.85% | ||||
| PartyScene | -0.27% | -0.61% | -1.03% | 2.92% | -0.17% | -0.36% | -0.16% | 2.67% | -0.17% | -0.27% | 0.03% | 2.29% | ||||
| RaceHorsesC | -0.95% | -1.23% | -0.39% | 5.75% | -0.23% | -0.13% | 0.30% | 3.55% | -0.37% | -0.40% | -0.06% | 2.72% | ||||
| Average | -1.26% | -1.43% | -1.41% | 7.45% | -0.60% | -0.58% | -0.37% | 5.09% | -0.50% | -0.46% | -0.29% | 3.58% | ||||
| ClassE (1280x720) | FourPeople | -0.93% | -0.64% | -0.17% | 3.56% | -0.46% | 0.29% | -0.36% | 2.47% | -0.41% | -0.35% | -0.21% | 1.37% | |||
| Johnny | -1.89% | -1.29% | -1.30% | 5.58% | -1.12% | -2.02% | 0.46% | 3.54% | -0.67% | -0.25% | -0.34% | 2.30% | ||||
| KristenAndSara | -0.68% | -1.63% | 0.02% | 3.35% | -0.38% | -0.80% | 0.66% | 2.19% | -0.48% | -0.77% | -0.18% | 1.36% | ||||
| Average | -1.17% | -1.19% | -0.48% | 4.16% | -0.66% | -0.84% | 0.25% | 2.73% | -0.52% | -0.46% | -0.24% | 1.68% | ||||
| Overall | -0.82% | -0.86% | -0.78% | 4.58% | -0.49% | -0.45% | -0.13% | 3.33% | -0.37% | -0.14% | -0.23% | 1.88% | ||||
| ClassD (416x240) | BasketballPass | -0.61% | -0.38% | -1.17% | 8.58% | -0.48% | -0.97% | -0.74% | 5.63% | -0.44% | 0.55% | -0.21% | 2.78% | |||
| BQSquare | -0.78% | -2.28% | 3.65% | 5.82% | -0.42% | -1.73% | 1.76% | 4.13% | -0.53% | -0.60% | -0.37% | 2.68% | ||||
| BlowingBubbles | -1.02% | -1.17% | -1.98% | 5.85% | -0.63% | -0.16% | 0.39% | 4.41% | -0.50% | -0.81% | -1.06% | 3.40% | ||||
| RaceHorses | -0.83% | -0.21% | -0.90% | 6.28% | -0.10% | 0.88% | 0.15% | 3.68% | -0.34% | 0.94% | 0.49% | 3.03% | ||||
| Average | -0.81% | -1.01% | -0.10% | 6.63% | -0.41% | -0.49% | 0.39% | 4.46% | -0.45% | 0.02% | -0.29% | 2.97% | ||||
VIII Syntax Design
Syntax elements of SAIP are coded into the bitstream using entropy coding. In the high-level syntax (HLS), two SAIP enabling flags (SAIP_Merge, SAIP_MMVD) and the maximum candidate number of the secondary candidate list are coded in the sequence parameter set. The enabling flags are coded with fixed-length code, and the maximum candidate number is coded with the 0-th order exponential-Golomb code.
At the coding unit level, the SAIP syntax elements are signaled inside the merge data syntax. Figure 8 shows the changes in the syntax tree of merge data before and after adding SAIP syntax elements. To minimize the impact of the newly increased syntax elements on coding the syntax of other modes, the original syntax tree of merge data has been redesigned. In the new syntax elements tree, saip_flag is signaled to indicate whether it is in SAIP mode or GPM (GEO) mode, and saip_merge_flag is signaled to indicate whether it is in SAIP-Merge mode or SAIP-MMVD mode.
Table I is the syntax elements table for SAIP. It contains four parts, including the syntax elements of the primary region (Merge: saip_merge_idx, MMVD: saip_mmvd_cand_idx, saip_mmvd_distance_idx and saip_mmvd_direction_idx), secondary region (saip_back_idx), mode flag (saip_flag and saip_merge_flag), and the reverse index (saip_reverse_idx). These elements are coded by using the context-based adaptive binary arithmetic coding (CABAC) engine. To code the newly increased syntax elements efficiently, the mode flag, reverse index, and the syntax elements of the secondary region are coded by using a new context model.
IX Experiments
IX-A Experimental Settings
IX-A1 Learned Segmentation Inference
@@
IX-A2 Evaluation Configurations
To evaluate the performance of SAIP in VVC, SAIP is implemented into the VVC reference software VTM-12.0 222https://vcgit.hhi.fraunhofer.de/jvet/VVCSoftware_VTM. Three main profile configurations (Low-delay P (LDP), Low-delay B (LDB), Random Access (RA)) and test sequences specified in [46] are used as the test conditions. Following the common test conditions (CTC)[46], four quantization parameters are tested: 22, 27, 32 and 37. For different quantization parameters, different segmentation configurations of the SCNet and STCN are chosen. For the measure of coding performance, the Bjontegaard Delta bit-rate (BD-rate) [47] is used as the objective metric to evaluate it. To further measure the characteristic of the proposed method on different sequences, the usage ratio () is used to validate the efficiency of mode selection, which is calculated by, , where indicates the number of coding units coded by the proposed method in P/B-frames, and indicates the total number of coding units in P/B-frames of the test sequences.
IX-B Performance
@@
| Class | Low-delay P | Low-delay B | Random Access | ||||||
| EncT | DecT | EncT | DecT | EncT | DecT | ||||
| A | - | - | - | - | 183% | 116% | |||
| B | 179% | 111% | 171% | 112% | 180% | 112% | |||
| C | 159% | 109% | 147% | 113% | 154% | 115% | |||
| E | 178% | 113% | 148% | 120% | 179% | 118% | |||
| Overall | 172% | 111% | 155% | 115% | 174% | 115% | |||
| D | 135% | 108% | 123% | 119% | 116% | 117% | |||
IX-B1 Overall Performance Under Common Test Conditions
The R-D performance of the entire segmentation-assisted inter prediction framework on the VVC common test sequences is illustrated in Table II . Y, U, and V represent the R-D performance gain of the three channels of YUV. We can see that our proposed SAIP can achieve on average, 0.82%, 0.49% and 0.37% (marked by font bold), and achieve up to 1.98%, 1.14% and 0.79% (marked by underline) BD-rate reduction (Y component) with high usage ratio on VTM-12.0 for all sequences under the LDP, LDB, and RA configurations. The experimental results show that the proposed framework performs better for sequences with more moving objects and complex motion fields, such as BasketballDrill, BQMall, CatRobot, and Marketplace.
In VVC reference software (VTM-12.0), the technology of inter prediction most relevant to SAIP is the geometric partitioning mode (GEO)[11]. Both ours and GEO aim to achieve a flexible description of the motion field and match the appropriate prediction for different regions in a block. In Table II , under the LDB and RA configurations, GEO is enabled in the anchor by default. In the comparison with anchor, SAIP achieves good performance with GEO enabled. It can be concluded that SAIP can efficiently deal with some complex motion scenarios that GEO is difficult to address.
IX-B2 Complexity
The encoding/decoding time complexity of our proposed method is illustrated in Table III . EncT and DecT represent the encoding/decoding time with the inference time of segmentation network, the network is inferred in parallel with the syntax parsing. The codec time is tested on CPU, and the model is Intel Core i7-11700 @2.50GHz. The inference time of network is tested on GPU, the models are aligned with the [31, 37], and the computational complexity is 119 kMAC/pixel.
| Class | Low-delay B | Random Access | |||||||||||||||||
| A1 | - | - | - | -0.20% | -5.69% | -0.20% | |||||||||||||
| A2 | - | - | - | -0.21% | -12.08% | -0.32% | |||||||||||||
| B | -0.31% | -1.17% | -0.30% | -0.19% | -6.96% | -0.26% | |||||||||||||
| C | -0.61% | -1.41% | -0.60% | -0.32% | -3.22% | -0.50% | |||||||||||||
| E | -0.88% | -1.29% | -0.66% | - | -8.35% | -0.52% | |||||||||||||
| Overall | -0.55% | -1.28% | -0.49% | -0.23% | -7.00% | -0.37% | |||||||||||||
| D | - | -1.02% | -0.41% | - | -0.83% | -0.45% | |||||||||||||
1
2
The performance of SAIP on top of VTM-12.0.
| Low-delay B Configuration (%) | ||||||||||||||||||||||||||||
@@ IX-B2 Complexity
| Name | Y | Y | U | V | Y | U | V | Y | U | V | Y | U | V | |||||||||||||||
| MarketPlace | - | - | - | - | -16.81% | -12.53% | -20.41% | -4.86% | 20.24% | 22.04% | -2.38% | -2.09% | -2.38% | |||||||||||||||
| RitualDance | - | - | - | - | -25.05% | -20.87% | -20.66% | -8.90% | 17.14% | 16.08% | -2.15% | -1.98% | -1.91% | |||||||||||||||
| Cactus | - | -1.00% | -0.90% | -1.00% | -20.36% | -6.45% | -6.35% | -9.88% | 22.29% | 22.37% | -0.80% | -0.77% | -0.86% | |||||||||||||||
| BasketballDrive | -0.08% | -0.70% | -0.80% | -1.10% | -24.02% | -21.09% | -21.25% | -9.47% | 27.71% | 23.75% | -2.20% | -3.00% | -2.88% | |||||||||||||||
| BQTerrace | - | -1.60% | -2.00% | -1.70% | -16.10% | -15.34% | -28.19% | -8.08% | 18.61% | 11.56% | -1.48% | -1.01% | -0.94% | |||||||||||||||
| Kimono | -0.14% | -0.90% | -1.00% | -1.40% | -15.01% | 4.40% | 0.68% | -6.15% | 20.14% | 20.14% | -1.17% | -1.00% | -0.87% | |||||||||||||||
| ParkScene | - | -0.60% | -0.40% | -0.70% | -16.60% | -3.55% | -4.14% | -6.89% | 20.26% | 22.02% | -0.23% | -0.63% | -0.55% | |||||||||||||||
| BasketballDrill | - | -1.30% | -1.50% | -1.30% | -29.14% | -15.11% | -14.63% | -19.76% | 9.37% | 10.45% | -4.34% | -4.27% | -4.91% | |||||||||||||||
| BQMall | - | -1.50% | -1.10% | -1.00% | -22.74% | -14.37% | -13.82% | -12.03% | 14.86% | 14.43% | -4.47% | -5.91% | -5.14% | |||||||||||||||
| PartyScene | - | -1.80% | -1.40% | -1.10% | -16.82% | 0.69% | -1.07% | -9.07% | 21.47% | 22.33% | -1.16% | -1.10% | -1.37% | |||||||||||||||
| RaceHorsesC | -0.95% | -0.60% | -0.90% | -0.80% | -12.52% | 10.91% | 10.28% | -5.06% | 30.48% | 34.81% | -2.21% | -2.97% | -3.63% | |||||||||||||||
| BasketballPass | - | -0.80% | -0.80% | -1.00% | -17.88% | -10.36% | -6.16% | -7.69% | 23.35% | 24.07% | -2.41% | -3.60% | -4.02% | |||||||||||||||
| BQSquare | - | -1.90% | -1.60% | -2.00% | -13.55% | -10.62% | -12.38% | -8.34% | 7.68% | 5.87% | -1.26% | 0.17% | 1.14% | |||||||||||||||
| BlowingBubbles | -0.30% | -1.70% | -1.40% | -1.60% | -12.62% | 6.62% | 2.90% | -5.55% | 24.08% | 24.18% | -2.18% | -2.35% | -1.14% | |||||||||||||||
| RaceHorses | - | -1.00% | -1.20% | -1.00% | -13.47% | 7.69% | 7.84% | -5.10% | 29.82% | 29.66% | -2.11% | -2.81% | -2.30% | |||||||||||||||
| FourPeople | - | -0.70% | -0.60% | -1.20% | -25.13% | -28.39% | -29.28% | -10.46% | 0.92% | -0.80% | -2.23% | -1.48% | -1.86% | |||||||||||||||
| Johnny | - | -0.90% | -1.00% | -1.20% | -25.81% | -30.38% | -34.35% | -10.25% | 9.35% | 6.11% | -3.80% | -3.04% | -2.73% | |||||||||||||||
| KristenAndSara | - | -1.10% | -1.30% | -0.90% | -23.97% | -26.32% | -28.35% | -8.89% | 6.64% | 5.79% | -1.53% | -1.07% | 2.21% | |||||||||||||||
| Average | -0.37% | -1.13% | -1.12% | -1.13% | -20.72% | -12.74% | -15.11% | -8.69% | 18.02% | 17.49% | -2.12% | -2.16% | -1.90% | |||||||||||||||
@@
| Low-delay P Configuration (%) | Random Access Configuration (%) | ||||||||||||||||||||||||||||||